OpenStudy (anonymous):

The ratio of Evan’s red toy cars to blue toy cars is 3 : 4. Evan has 12 red toy cars. How many blue toy cars does Evan have?

OpenStudy (anonymous):

9 is the right answer because 3/4ths of 12 is nine.

OpenStudy (anonymous):

3*3=9,3*4=12

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Oh I just re-read the question before you typed that, and I was about to say wait I misread it! lol. Glad you got it right!

OpenStudy (anonymous):

He has three fourths the amount of red cars to blue cars so yes it s sixteen because 12 is three fourths of 16. have a nice day!

OpenStudy (anonymous):

Thanks for the medal!

OpenStudy (anonymous):

In a bag of red and yellow cookies, the ratio of red cookies to yellow cookies is 4:5. If the bag contains 80 yellow cookies, how many red cookies are there? A) 56 B) 60 C) 64 D) 68

OpenStudy (anonymous):

it would be 64. 64/4=16 and 80/5=16 do you understand?

OpenStudy (anonymous):

64/80 is equal to 4/5 so the answer is 64